OpenStudy (rainbow_dashie):

What is the solution to the equation \[8-2b=-\frac{ 2 }{ 3 }(12b+6)\] ?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you know how to distribute on the right hand side of the equal sign?

OpenStudy (rainbow_dashie):

no not really

OpenStudy (misty1212):

\[-\frac{2}{3}\times 12=?\]

OpenStudy (rainbow_dashie):

-8

OpenStudy (misty1212):

ok good so we got \(-8b\) now how about \[-\frac{2}{3}\times 6\]?

OpenStudy (rainbow_dashie):

-4

OpenStudy (misty1212):

ok so now we have \[8-2b=-8b-4\]

OpenStudy (anonymous):

im still here different account

OpenStudy (misty1212):

your final job is to solve that for \(b\) as usual oh did you change your picture?

OpenStudy (anonymous):

no i have two account that i like to go on and off on.

OpenStudy (anonymous):

and how do i do that

OpenStudy (misty1212):

oh are you still from england?

OpenStudy (anonymous):

ya.

OpenStudy (misty1212):

and still 13 and taken?

OpenStudy (anonymous):

lol ay

OpenStudy (misty1212):

\[8-2b=-8b-4\]add \(8b\) get \[8+6b=-4\] subtract \(8\) get \[6b=-12\] and finally divide by \(6\)

OpenStudy (anonymous):

-0.5

OpenStudy (misty1212):

if you are from england, why do you go to georgia cyber academy instead of a real school?

OpenStudy (misty1212):

oh no darlling \(\frac{-12}{6}=-2\) not \(-.5\)!

OpenStudy (anonymous):

because i lived in the usa for 2 and a half years and they kept me in for another year

OpenStudy (anonymous):

Oh sorry XD

OpenStudy (misty1212):

oh too bad ... in any case answer is \(-2\) ok?

OpenStudy (misty1212):

don't they have real schools in georgia?

OpenStudy (anonymous):

ya but when i was in there they didn't have good schools, cussing and stolen stuff. and i ment they kept me n there for the rest of the year. sorry for confusion XD

OpenStudy (misty1212):

hey no problem get out of georgia and back to england as soon as you can!!

OpenStudy (anonymous):

\[\frac{ -2n-2 }{ 3 } = 12\]

OpenStudy (anonymous):

one more? OuO